there are 10 pints of a 40% acid solution that is being added to 4 pints of an 80% solution. what will be the level of acid concentration in the mixture?
Answer by stanbon(75887)

there are 10 pints of a 40% acid solution that is being added to 4 pints of an 80% solution. what will be the level of acid concentration in the mixture?
----
Equation:
acid + acid = acid
0.40*10 + 0.80*4 = x*(14)
----
Multiply thru by 100 to get:
40*10 + 80*4 = 1400x
40 + 320 = 1400x
x = 360/1400
x = 0.257 = 25.7%
=========================
Cheers,
Stan H.
=========================